# Uni-NaVid: A Video-based Vision-Language-Action Model for Unifying Embodied Navigation Tasks

### v1

@@ 4.3 Ablation Study
We present a visualization of the training strategy’s performance in Figure 5. This comparison contrasts training on a single navigation task with training across multiple tasks. The results (Figure 5 (a)) illustrate the synergistic benefits of multi-task learning, yielding consistent performance gains across all navigation tasks. Notably, VLN, ObjectNav, and EQA exhibit more substantial improvements, while Following shows comparatively smaller gains. We attribute this to the lower dependence of the following task on historical context. Additionally, we investigate the influence of data scale on performance (Figure 5 (b)), observing that performance improves across all navigation tasks with larger data volumes. However, the incremental gain diminishes (from 3M to 6M samples), potentially due to limitations in the data diversity of simulators.
@@
We conduct experiments to evaluate the effectiveness of the training strategy and visual token designs (Tab. 9). Our results indicate that the absence of <NAV> and VQA data leads to a performance decline across all tasks. Notably, the performance drop is most pronounced in EQA, as the lack of <NAV> special token makes the model misinterpret whether it should answer questions or output actions. Additionally, without VQA data, the agent’s ability to answer questions using visual information is greatly impaired.
@@
From the performance of different memory designs, we find that both short-term and long-term memory visual tokens contribute to performance improvements. In particular, the VLN task shows the most significant performance drop ($-80.3\%$ SR) when visual memory is removed, as the lack of memory hinders the alignment of visual history with instructions. For the Following task, the absence of memory results in only a minor performance decline ($-8\%$ SR), as this task primarily relies on recent frames to track the target. Additional ablation studies are provided in the Supplementary Material.
@@
Table 9: Ablation study on both training strategy and architecture.
5 Conclusion
We introduce a vision-language-action model, Uni-NaVid, designed to acquire general embodied navigation skills through multi-task navigation data. To efficiently encode the extended video sequences encountered in navigation, we develop an online visual token merging strategy that separately processes current observations, short-term observations, and long-term observations. This design allows our approach to operate at an average speed of 5 Hz. Extensive experiments across publicly available benchmarks reveal our model’s superior performance. The ablation studies demonstrate the synergy achieved by learning across multiple navigation tasks and prove the effectiveness of our key designs. In the future, we will explore the synergy between more fundamental embodied skills, integrating both manipulation and navigation.

### v2

@@ VI-D Ablation Study
Visualization of training strategy. We present a visualization of the training strategy’s performance in Figure 8. In Fig. 8 (a), we compare training on a single navigation task with training across multiple tasks. The results demonstrate the synergistic benefits of multi-task learning, which yields consistent performance improvements across all navigation tasks. Notably, VLN, ObjectNav, and EQA exhibit more significant improvements, while Following shows relatively smaller gains. We attribute this difference to the lower reliance of the Following task on historical context. Additionally, we investigate the influence of data scale on navigation performance (Figure 8 (b)). We observe that performance improves across all navigation tasks with larger data volumes. However, the incremental gain diminishes (from 3M to 6M samples), potentially due to limitations in the data diversity of simulators. Specifically, for the Following task, the reason for the slower convergence is the heavy occlusion caused by obstacles or other humans. This highlights the need for more high-quality following data samples, which can enable our model to learn more effectively and perform better in highly dynamic environments.
@@
Ablation on training strategy and architecture. We conduct experiments to evaluate the effectiveness of the training strategy and token merging designs (Tab. X). Our results indicate that the absence of <NAV> and VQA data leads to a performance decline across all tasks, similar findings can be found in [14, 111]. Notably, the performance drop is most obviously in EQA, as the lack of <NAV> special token makes the model misinterpret whether it should answer questions or output actions. Additionally, without VQA data, the agent’s ability to answer questions drops significantly, almost rendering it incapable of correctly answering questions. We believe this is due to the catastrophic forgetting problem in LLMs, where the model loses open-world knowledge by being trained solely on navigation-related data.
@@
From the performance of different memory designs, we find that both short-term and long-term memory visual tokens contribute to performance improvements. In particular, the VLN task shows the most significant performance drop ($-80.3\%$ SR) when visual memory is removed, as the lack of memory hinders the alignment of visual history with instructions. For the Following task, the absence of memory results in only a minor performance decline ($-8\%$ SR), as this task primarily relies on recent frames to track the target. Additional ablation studies on architecture and hyperparameters are provided in the Supplementary Material.
@@
TABLE X: Ablation study on training strategy and architecture. For each ablation type, we retrain the entire model and evaluate its performance across four navigation tasks.
VII Limitations
Despite the promising results, Uni-NaVid has several limitations. First, Uni-NaVid is trained and evaluated on four well-defined navigation tasks, while there exists a large body of literature on insightful and practical navigation datasets [86, 112]. We believe that collecting data from these datasets could further enhance the navigation capabilities of our method. Second, our method is designed to acquire multi-task navigation capabilities under the assumption that the robot is of standard size (see Section V-A). To extend it to robots of general sizes, a convincing approach is to incorporate prior knowledge of the robot’s size, as demonstrated in [24, 29]. Third, our method is currently limited to predicting simple trajectories composed of a short horizon of future low-level discrete actions. This limitation could be alleviated by extending the moel to predict continuous and smooth trajectories with techniques from motion planning [79, 81] or autonomous driving [18, 52].
VIII Discussion and Conclusion
In this paper, we introduce an efficient vision-language-action (VLA) model, Uni-NaVid, designed to acquire general embodied navigation skills through learning multi-task navigation data. To efficiently encode the online-captured video sequences during navigation, we develop an online visual token merging mechanism that separately processes current observations, short-term observations, and long-term observations. This design enables our approach to operate at an average speed of 5 Hz. We also collect 3.6 million navigation data points across four highly demanded embodied navigation tasks, including vision-and-language navigation, object goal navigation, embodied question answering, and human following. Extensive experiments and ablation studies demonstrate that our method achieves SOTA-level performance using only monocular videos as input, highlighting our model’s superior capability in learning multiple navigation tasks. Moreover, we deploy Uni-NaVid in real-world environments, demonstrating impressive generalizability and versatile navigation performance in real worlds.
Future works. Our work serves merely as a starting point of general-purpose navigation, and we hope it will inspire future directions in this field:
  • Benchmarking. With the consistent development of embodied navigation, there is a growing need for general-purpose navigation benchmarking. Such a benchmark would help researchers better position their work and drive progress in the navigation community.
  • Architecture. We would like to further enhance the practicality of our architecture by tackling very long-horizon tasks (e.g., navigating across buildings) and incorporating advanced motion planning techniques [81, 18].
  • Application. We would like to apply our method to applications such as robotic guide dogs and home service robots. Additionally, we are excited to extend this technique to other embodied AI tasks, such as mobile manipulation [110, 92, 55].